OpenStudy (anonymous):

While taking a walk along the road where you live, you accidentally drop your glove, but you don't know where.

OpenStudy (vishal_kothari):

what is this?

OpenStudy (lilai3):

ikr...wut is this.. and where did u drop it...or sumthing?

OpenStudy (anonymous):

While taking a walk along the road where you live, you accidentally drop your glove, but you don't know where. The probability density p(x) for having dropped the glove x kilometers from home (along the road) is p(x)=3e^(−3x ) for x≥0 a.What is the probability that you dropped it within 1 kilometer of home? b.At what distance y from home is the probability that you dropped it within y km of home equal to 0.95? thanks!!

OpenStudy (zarkon):

a)\[\int\limits_{0}^{1}p(x)dx\] b) solve \(\displaystyle\int\limits_{0}^{y}p(x)dx=.95\) for y

OpenStudy (zarkon):

\[\displaystyle\int\limits_{0}^{y}p(x)dx=1-e^{-3y}\] \solve \[1-e^{-3y}=.95\] \[1-.95=e^{-3y}\] \[.05=e^{-3y}\] \[\ln(.05)=-3y\] \[y=\frac{\ln(.05)}{-3}\approx.998577\]

OpenStudy (anonymous):

thanks!